Status Report

Hubble Space Telescope Daily Report #2705 09/12/00

By SpaceRef Editor
September 12, 2000
Filed under

PERIOD COVERED:  0000Z (UTC) 09/11/00 – 0000Z (UTC) 09/12/00
Daily Status Report as of 256/0000Z
    1.1 Completed Five Sets of WF/PC-2 8826 (Cycle 9 Supplemental Darks pt 1)
        The WF/PC-2 was used to perform a T dark calibration program that obtains three dark frames every day in order to provide data for monitoring and characterizing the evolution of hot pixels.  There were no reported anomalies.
    1.2 Completed Four Sets of STIS/CCD 8837 (CCD Dark Monitor-Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
    1.3 Completed WF/PC-2 8701 (Multiplicity among Very-Low Mass Stars and Brown Dwarfs in Alpha Persei and the Pleiades)
        The WF/PC-2 was used to observe open clusters which provide excellent hunting grounds for brown dwarf {BD} searches.  The AlphaPer and Pleiades clusters are young, nearby and have low extinction.  They are cosidered to be the best-suited places to study the Substellar Mass.  A dozen of cool faint AlphaPer and Pleiades members have been confirmed as “bona-fide” BDs with the lithium test.  By comparison with those, about 60 very good cluster BD candiates are currently known.  The Pleiades SMF obtained using the most recent deep large CCD surveys indicates that BDs are quite numerous but do not make a significant contribution to the total cluster mass.  The observations completed with no reported problems.
    1.4 Completed STIS/CCD 8808 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation III)
        The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted, parallel observations as part of a POMS test proposal.  The observations completed with no anomalous activity.
    1.5 Completed WF/PC-2 8601 (A Snapshot Survey of Probable Nearby Galaxies)
        The WF/PC-2 was used to continue the very successful snapshot survey in order to use the high spatial resolution of HST to determine whether selected galaxies are nearby on the basis of resolution into stars, and the magnitudes and colors of the brightest stars.  There were no reported problems.
    1.6 Completed STIS/CCD 8260 (Searching for the Hydrogen Reionization Edge of the Universe at 5HST)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe one of 36 parallel orbits {4–5 fields of 5–8 orbits each} to constrain the H Lyman-edge in emission that marks the transition from a neutral to a fully ionized IGM at a predicted zion~eq5–15.  This edge is due to recombination from the H Lyman series and Lyman continuum, and can be used to constrain zion, one of the most important unknown quantities in large scale structure and cosmology.  The proposal completed nominally.
    1.7 Completed Six Sets of WF/PC-2 8811 (Cycle 9 Standard Darks)
        The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels.  Over an extended period these data will also provide a monitor of radiation damage to the CCDs.  The proposal completed with no reported problems.
    1.8 Completed WF/PC-2 8812 (Cycle 9 Internal Monitor)
        The WF/PC-2 was used to obtain routine internal monitors for WFPC2 in order to monitor the health of the cameras.  A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows.  There were no reported problems.
    1.9 Completed STIS/CCD/MA1 8190 Luminous Infrared Galaxies- Superstarbursts or AGN)
        The Space Telescope Imaging Spectrograph (CCD and MA1) was used to obtain high resolution spectroscopy of four of the nearest `warm’ ultraluminous infrared galaxies (ULIGs) that were discovered in the IRAS database, and for which we have recently obtained WFPC2 images in B, I filters.  These objects are thought to represent a critical evolutionary link between ULIGs and optically selected QSOs.  The observations completed with no reported problems.
    1.10 Completed Four Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation)
        The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program.  The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working.  The proposal completed with no reported problems.
    1.11 Completed STIS/CCD 8700 (A STIS Spectroscopic Snapshot Survey of 3CR Radio Galaxies : The Nature of the Unresolved Nuclei)
        The Space Telescope Imaging Spectrograph (CCD) was used to perform STIS snapshot long-slit spectroscopy of these nuclei over 2900- 10300 AA to classify them {QSO, NLRG, BLRG, BL Lac, ldots} and establish their AGN characteristics down to the lowest nuclear luminosities, while still within powerful 3C radio galaxies.  The primary goals are : {1} to test unification schemes by identifying previously undetected signatures of low-level AGN/QSO activity in their spectra in the form of weak, broad emission lines, featureless continua, and UV excess, {2} to constrain the FR I/BL Lac connection by measuring the relative contribution of thermal {starburst}, non-thermal {optical synchrotron}, and line emission, and {3} to perform line and continuum diagnostics to determine the dominant ionization mechanisms in the circumnuclear regions {tenths of arcseconds} and the role of dust extinction in the NLRGs and BLRGs in the form of geometrically thick or thin tori and foreground dust.  The proposal completed as planned, with no reported anomalies.
    1.12 Completed STIS/CCD/MA1 8103 (A Global Picture of White Dwarfs in Cataclysmic Variables)
        The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of the white dwarf star that is located in the cataclysmic variable star system known as CP-ERI.  The observations were executed as scheduled, and no problems were reported
    1.13 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
    1.14 Completed WF/PC-2/STIS/CCD 8191 (The Faintest Radio Galaxies: Interacting Starbursts at z<1)
        Deep WF/PC-2 imaging and Space Telescope Imaging Spectrograph (CCD) slitless spectroscopy of two of our ultra-deep VLA fields were
performed.  There were no reported problems.
    1.15 Completed WF/PC-2/STIS/CCD 8573 (Newborn Planets and Brown Dwarf Companions in IC 348)
        The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to search for young giant planets and brown dwarfs around ~100 low-mass stars and brown dwarfs in the nearby cluster IC 348.  The observations completed with no reported problems.
    1.16 Completed WF/PC-2 8720 (Masses and Multiplicity of Nearby Free- floating Methane and L Dwarfs)
        The WF/PC-2 was used to observe 50 very-low-mass objects in the solar neighborhood with spectral types of L0 and later {including several dwarfs with Methane absorption bands in their atmospheres}.  Tthese objects will be observed in two filter bands with the aim to identify close companions, measure their colors, and to obtain first epoch data of the newly discovered binaries.  The observations completed nominally.
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions: 8
Successful: 8
  Scheduled Re-acquisitions: 8
Successful: 8
    2.2 FHST Updates:
Scheduled: 19
Successful: 19
    2.3 Operations Notes:
        Using ROP SR-1A, the SSR EDAC error counter was cleared twice.
        As documented in HSTAR 7846, telemetry parameter SMA1V1PG was out-of-limit for one telemetry sample with a value of 0.0285 at 255/204701Z.
        SSA transmitter #1 was turned on at 256/0031Z and turned off at 256/0031Z, per ROP IC-2.
        The ephemeris uplink was performed at 256/0056Z as directed by ROP DF-07B.
        Continuation of normal science observations and calibrations.

SpaceRef staff editor.